2 次 (拡大) 体の 4-rank に関する文献 (2012 年 2 月更新版)

[1] L. Rédei : A másodfokú képzetes számtest osztály-számáról, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 44 (1927), 230– 246; JFM 53.0132.04 (reviewed by Koenig, D.) [html, DVI, ps, PDF, TeX].

[2] L. Rédei : Über die Klassenzahl des imaginären quadratischen Zahlkörpers, J. Reine Angew. Math. 159 (1928), Heft 4, 210–219; doi: 10.1515/crll.1928.159.210 [PDF (1367 KB)];

PPN243919689_0159_log22.pdf (1,03 MB), GDZ; [PDF],

DigiZeitschriften; JFM 54.0178.07 (reviewed by Rothe, H.) [html, DVI, ps, PDF, TeX].

[3] L. Rédei : A másodfokú valós számtest osztályszámáról s alapegységéről, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 48 (1931), 648–682; Zbl 0005.09803 (reviewed by Kalmár) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften.

[4] L. Rédei : A másodfokú számtest osztály-számáról, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 48 (1931), 683–707; Zbl 0005.09901 (reviewed by Kalmár) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften.

[5] L. Rédei : A másodfokú számtest osztálycsoportjának 4gyel osztható invariánsai, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 49 (1932), 338–363; Zbl 0008.19502 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 59.0941.02 (reviewed by Zányi, L.) [html, DVI, ps, PDF, TeX].

[6] L. Rédei : A másodfokú számtest osztálycsoportjának 8cal osztható invariánsai, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 50 (1933), 195–218; Zbl 0009.29401 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 60.0934.01 (reviewed by Zányi, L.) [html, DVI, ps, PDF, TeX].

[7] L. Rédei : A másodfokú számtest egyik tételénak uj bizonyitása, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 50 (1933), 219–230; Zbl 0009.29301 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 60.0934.02 (reviewed by Zányi, L.) [html, DVI, ps, PDF, TeX].

[8] L. Rédei und H. Reichardt : Die Anzahl der durch 4 teilbaren Invarianten der Klassengruppe eines beiebigen quadratischen Zahlkörpers, J. Reine Angew. Math. 170 (1934), Heft 2, 69–74; doi: 10.1515/crll.1934.170.69 [PDF (1066 KB)];

PPN243919689_0170_log7.pdf (870,16 KB), GDZ; [PDF],

DigiZeitschriften; Zbl 0007.39602 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 59.0192.01 (reviewed by Scholz, A.) [html, DVI, ps, PDF, TeX];

[9] H. Reichardt : Zur Struktur der absoluten Idealklassengruppe im quadratischen Zahlkörper, J. Reine Angew. Math. 170 (1934), Heft 2, 75–82; doi: 10.1515/crll.1934.170.75

[PDF (1373 KB)]; PPN243919689_0170_log8.pdf (1,09 MB),

GDZ; [PDF], DigiZeitschriften; Zbl 0007.39603 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 59.0192.02 (reviewed by Scholz, A.) [html, DVI, ps, PDF, TeX].

[10] L. Rédei : Felső korlát a másodfokú számtest abszolút osztálycsoportjának 4-gyel osztható invariánsai számára, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 51 (1934), 219– 226; Zbl 0010.33801 (reviewed by O. Szász) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 60.0934.03 (reviewed by Zányi, L.) [html, DVI, ps, PDF, TeX].

[11] L. Rédei : A másodfokú számtest alapegységéről és az abszolút osztálycsoport 8-cal osztható invariánsairól, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 51 (1934), 227–259; Zbl 0010.33802 (reviewed by O. Szász) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 60.0934.04 (reviewed by Zányi, L.) [html, DVI, ps, PDF, TeX].

[12] L. Rédei : Arithmetischer Beweis der Satzes über die Anzahl der durch vier teilbaren Invarianten der absoluten Klassengruppe im quadratischen Zahlkörper, J. Reine Angew. Math. 171 (1934), Heft 1, 55–60; doi: 10.1515/crll.1934.171.55

[PDF (1025 KB)]; PPN243919689_0171_log9.pdf (817,72

KB), GDZ; [PDF], DigiZeitschriften; Zbl 0009.05101 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 60.0125.02 (reviewed by Hasse, H.) [html, DVI, ps, PDF, TeX].

[13] L. Rédei : Eine obere Schranke der Anzahl der durch vier teilbaren Invarianten der absoluten Klassengruppe im quadratischen Zahlkörper, J. Reine Angew. Math. 171 (1934), 61–64; doi: 10.1515/crll.1934.171.61 [PDF (676 KB)];

PPN243919689_0171_log10.pdf (482,05 KB), GDZ; [PDF],

DigiZeitschriften; JFM 60.0125.03 (reviewed by Hasse, H.) [html, DVI, ps, PDF, TeX].

[14] L. Rédei : Über die Grundeinheit und die durch 8 teilbaren Invarianten der absoluten Klassengruppe im quadratischen Zahlkörper, J. Reine Angew. Math. 171 (1934), Heft 3, 131–148; doi: 10.1515/crll.1934.171.131 [PDF (2323 KB)];

PPN243919689_0171_log14.pdf (1,99 MB), GDZ; [PDF],

DigiZeitschriften; Zbl 0009.29302 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 60.0125.04 (reviewed by Hasse, H.) [html, DVI, ps, PDF, TeX]; Berichtigung in [16], S. 55.

[15] L. Rédei : Néhány középértékkérdésről másodfokú számtestekben, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 54 (1935), 45–116; Zbl 0014.05201 (reviewed by O. Szász) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 62.0167.01 (reviewed by Lipka, S.) [html, DVI, ps, PDF, TeX].

[16] L. Rédei : Néhány újabb kongruenciafeltétel a másodfokú számtest abszolút osztálycsoportjának nyolocal osztható invariánsaira, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 54 (1935), 736–768; Zbl 0015.29303 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 62.1155.03 (reviewed by Lipka, S.) [html, DVI, ps, PDF, TeX].

[17] L. Rédei : Über einige Mittelwertfragen im quadratischen Zahlkörper, J. Reine Angew. Math. 174 (1936), Heft 1/2, 15–55; doi: 10.1515/crll.1936.174.15 [PDF (5476 KB)];

PPN243919689_0174_log5.pdf (4,93 MB), GDZ; [PDF],

DigiZeitschriften; Zbl 0013.00304 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 61.0171.02 (reviewed by Reichardt, H.) [html, DVI, ps, PDF, TeX].

[18] L. Rédei : Egy új számelméleti jel, alkalmazással a másodfokú számtestek elméletére I, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 56 (1937), 807–847; Zbl 0019.00401 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 63.0915.01 (reviewed by Lipka, S.) [html, DVI, ps, PDF, TeX].

[19] L. Rédei : Másodfokú számtestek absolút osztály-csoportjának 2-, 4- és 8-cal osztható invariánsai, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 56 (1937), 848–853; Zbl 0019.00501 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 63.0916.01 (reviewed by Lipka, S.) [html, DVI, ps, PDF, TeX].

[20] L. Rédei : Másodfokú számtestek absolút osztály-csoportjának 4-gyel osztható invariánsainak számosságára vonatkozó néhány középértékkérdés, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 57 (1938), 88–104; Zbl 0018.38901 (reviewed by O. Szász) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 64.0990.02 (reviewed by Schmid, H. L.) [html, DVI, ps, PDF, TeX].

[21] L. Rédei : Egy új számelméleti jel, alkalmazással a másodfokú számtestek elméletére II, Math. Naturwiss. Anz. Ungar. Akad. Wiss. [html, PDF, DVI, PostScript, BibTeX]; Zbl 0020.10104 (reviewed by Reichardt) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 64.0990.03 (reviewed by Schmid, H. L.) [html, DVI, ps, PDF, TeX].

[22] L. Rédei : Ein neues zahlentheoretisches Symbol mit Anwendungen auf die Theorie der quadratischen Zahlkörper. I, J. Reine Angew. Math. 180 (1939), Heft 1/2, 1–43; doi: 10.1515/crll.1939.180.1 [PDF (5642 KB)];

PPN243919689_0180_log4.pdf (5,08 MB), GDZ; [PDF], Di-

giZeitschriften; Zbl 0021.00701 (reviewed by Reichardt) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 65.0106.03 (reviewed by Reichardt, H.) [html, DVI, ps, PDF, TeX].

[23] L. Rédei : Die diophantische Gleichung mx^{2} + ny^{2} = z^{4}, Monatsh. Math. Phys. 48 (1939), 43–60; DOI: 10.1007/BF01696163 [l2982m524vpr22j3.pdf (756.0 KB)]; 57 (1938), 488–500; MR0000829 (1, 134f) (reviewed by A. Brauer; reprinted in Reviews in Number Theory, volume 2, D32-1) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0021.29502 (reviewed by H. Reichardt) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 65.0133.03 (reviewed by Reichardt, H.) [html, DVI, ps, PDF, TeX].

[24] L. Rédei : Másodfokú számtestek gyűrűosztály-csoportjának páros részérűol, a Pell-féle és az rx^{2} + sy^{2} = z^{2n} diophantosi egyenletűol, I, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 62 (1943), 13–34; MR0023299 (9, 336h); (reviewed by A. Rényi; reprinted in Reviews in Number Theory, volume 5, R14-9) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0063.06449 (not reviewed).

[25] L. Rédei : Másodfokú számtestek gyűrűosztály-csoportjának páros részérűol, a Pell-féle és az rx^{2} + sy^{2} = z^{2n} diophantosi egyenletűol, II, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 62 (1943), 35–47; MR0023300 (9, 336i); (reviewed by A. Rényi; reprinted in Reviews in Number Theory, volume 5, R14-9) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0063.06449 (not reviewed).

[26] L. Rédei : Másodfokú számtestek gyűrűosztály-csoportjának páros részérűol, a Pell-féle és az rx^{2} + sy^{2} = z^{2n} diophantosi egyenletűol, III, Math. Naturwiss. Anz. Ungar. Akad. Wiss. 62 (1943), 48–62; MR0023301 (9, 336j); (reviewed by A. Rényi; reprinted in Reviews in Number Theory, volume 5, R14-9) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0063.06449 (not reviewed).

[27] L. Rédei : Die 2-Ringklassengruppe des quadratischen Zahlkörpers und die Theorie der Pellschen Gleichung, Acta Math. Acad. Sci. Hungar. 4 (1953), no. 1-2, 31–87; DOI: 10.1007/BF02020351 [g36r22857567471m.pdf (2.8 MB)]; MR0065589 (16, 450e) (reviewed by J. Tate; reprinted in Reviews in Number Theory, volume 5, R14-28) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0051.27101 (reviewed by H. Hasse) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften.

[28] D. Pamplün : Über die Klassenzahl und die Grundeinheit des reele-quadratischen Zahlkörpers, J. Reine Angew.

Math. 230 (1968), 167–210; doi: 10.1515/crll.1968.230.167

[PDF (4525 KB)]; PPN243919689_0230_log12.pdf (4,01

MB), GDZ; [PDF], DigiZeitschriften; MR0224590 (37 #189) (reviewed by O. H. Körner; reprinted in Reviews in Number Theory, volume 5, R14-68) [html, PDF, DVI, PostScript, Bib-TeX]; Zbl 0162.06901 (reviewed by by A. Yokoyama) [scanned image].

[29] P. Damey : Extensions cycliques de degré 4, non ramifieés d’un corps quadratiques sur Q, Séminaires Théorie de Nombres, 1968-1969, Univ. Bordeaux, Exp. No. 12, 1969, 6 pp.; Zbl 0287.12016 (reviewed by W. Zink) [scanned image].

[30] P. Damey et J.-J. Payan : Existence et construction des extensions galoisiennes et non abéliennes de degré 8 d’un corps de caractéristique différente de 2, J. Reine Angew. Math. 244 (1970), 37–54; doi: 10.1515/crll.1970.244.37 [PDF

(2338 KB)]; PPN243919689_0244_log6.pdf (2,01 MB), GDZ;

[PDF], DigiZeitschriften; MR0280466 (43 #6186) (reviewed by W.-D. Geyer; reprinted in Reviews in Number Theory, volume 5, R14-102) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0206.34401 (Autorreferat) [scanned image].

[31] L. Bouvier : Sur le 2-groupe des classes au sens restreint de certaines extensions biquadratiques de Q, C. R. Acad. Sci. Paris Sér. A-B 272 (1971), A193–A196; [PDF, JPEG, txt], Gallica; MR0277506 (43 #3239) (reviewed by J. Browkin; reprinted in Reviews in Number Theory, volume 5, R16-56) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0209.35701 (Autorreferat) [scanned image].

[32] L. Bouvier : Table des 2-rang, 4-rang et 8-rang du 2groupe des classes d’idéaux au sens restreint de Q(\sqrt{m}), m étant un entier relatif sans facteur carré tel que 1 < |m| < 10, 000, Enseignment. Math. (2) 18 (1972), 37–45;

ensmat-001_1972_18_1_a_002_d.pdf, SEALS; MR0337890

(49 #2659) (reviewed by O. Taussky-Todd; reprinted in Reviews in Number Theory, volume 5A, Y40-207) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0238.12005 (reviewed by W. Zink) [scanned image].

[33] G. Gras : Sur les l-classes d’idéaux dan les extensions cycliques relatives de degré premier l, Ann. Inst. Fourier (Grenoble) 23 (1973), Fasc. 4, 1–44; doi: 10.5802/aif.480

[AIF_1973__23_4_1_0.pdf]; AIF_1973__23_4_1_0.  djvu,
AIF_1973__23_4_1_0.pdf, Numdam; MR0360519 (50

#12967) (reviewed by J. Browkin; reprinted in Reviews in Number Theory, volume 5A, R26-203F) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0262.12007 (Autorreferat) [scanned image].

[34] P. Kaplan : Comparaison des 2-groupes des classes d’idéaux au sens large et au sens étroit d’un corps quadratique réel, Proc. Japan Acad. [日本學士院紀要] 50 (1974), No. 9, 688– 693; doi:10.3792/pja/1195518789, euclid.pja/1195518789 [PDF (568 KB)]; JOI:JST.Journal archive/pjab1945/50.688 [50_688.pdf (531 K)], Journal@rchive, JST; MR0384753 (52 #5626) (reviewed by P. Hartung; reprinted in Reviews in Number Theory, volume 5A, R14-260) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0337.12002 (reviewed by R. Bölling) [scanned image].

[35] P. Kaplan : Sur le 2-groupe des classes des corps quadratiques, Journées arithmétiques (Grenoble, 1973), Bull. Soc. Math. France Mem. 37 (1974), 115–116;

MSMF_1974__37__115_0.djvu, MSMF_1974__37__115_0.pdf,

Numdam; MR0399045 (53 #2896) (abstract) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0307.12003 (reviewed by R. Bölling) [scanned image].

[36] P. Kaplan : Sur le 2-groupe de classes d’idéaux des corps quadratiques, J. Reine Angew. Math. 283/284 (1976), 313– 363; doi: 10.1515/crll.1976.283-284.313 [PDF (4998 KB)];

PPN243919689_0283_0284_log29.pdf (4,44 MB), GDZ;

[PDF], DigiZeitschriften; MR0404206 (53 #8009) (reviewed by J. Martinet; reprinted in Reviews in Number Theory, volume 5A, R14-224G) [html, PDF, DVI, PostScript, BibTeX].

[37] H. Kisilevsky : The Rédei-Reichardt theorem—a new proof, Selected topics on ternary forms and norms (Sem. Number Theory, California Inst. Tech., Pasadena, Calif., 1974/75), Paper No. 6, California Inst. Tech., Pasadena, Calif., 1976, 4 pp.; MR0437494 (55 #10422) (reviewed by Y. Furuta; reprinted in Reviews in Number Theory, volume 5A, R14-289) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0369.12004 (reviewed by M. E. Keating) [scanned image].

[38] J. Carroll : The Rédei-Reichardt theorem, Selected topics on ternary forms and norms (Sem. Number Theory, California Inst. Tech., Pasadena, Calif., 1974/75), Paper No. 7, California Inst. Tech., Pasadena, Calif., 1976, 7 pp.; MR0437495 (55 #10423) (reviewed by Y. Furuta; reprinted in Reviews in Number Theory, volume 5A, R14-290) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0369.12005 (reviewed by M. E. Keating) [scanned image].

[39] J. Bloom : On the 4-rank of the strict class group of a quadratic number field. From L. Rédei, Arithmetischen Beweis der Satzes über die Anzahl der durch vier teilbaren Invarianten der absoluten Klassengruppe im quadratischen Zahlkörper (J. Reine Angew. Math. 171 (1934), 55–60), Selected topics on ternary forms and norms (Sem. Number Theory, California Inst. Tech., Pasadena, Calif., 1974/75), Paper No. 8, California Inst. Tech., Pasadena, Calif., 1976, 4 pp.; MR0437492 (55 #10420) (reviewed by Y. Furuta; reprinted in Reviews in Number Theory, volume 5A, R14-287) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0369.12006 (reviewed by M. E. Keating) [scanned image].

[40] B. Oriat : Relation entre les 2-groupes des classes d’ideaux au sens ordinaire et restreint de certains corps de nombres, Bull. Soc. Math. France 104

(1976), no. 3, 301–307; BSMF_1976__104__301_0.djvu,
BSMF_1976__104__301_0.pdf, Numdam; MR0435038 (55

#8000) (reviewed by P. Damey; reprinted in Reviews in Number Theory, volume 5A, R24-223) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0352.12007 (reviewed by H. Yokoi) [scanned image].

[41] B. Oriat : Relation entre les 2-groupes des classes d’idéaux des extensions quadratiques k(\sqrt{d}) et k(\sqrt{−d}), Journées Arithmétiques de Caen (Univ. Caen, Caen, 1976), Astérisque No. 41–42, Soc. Math. France, Paris, 1977, pp. 247–249; MR0506162 (58 #22013) (reviewed by Editors; reprinted in Reviews in Number Theory, volume 5A, R14-299F) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0348.12007 (résumé) [scanned image].

[42] B. Oriat : Relation entre les 2-groupes des classes d’idéaux des extensions quadratiques k(\sqrt{d}) et k(\sqrt{−d}), Ann. Inst. Fourier (Grenoble) 27 (1977), Fasc. 2, vii, 37–59; doi: 10.5802/aif.650 [AIF_1977__27_2_37_0.pdf];

AIF_1977__27_2_37_0.djvu, AIF_1977__27_2_37_0.pdf,

Numdam; MR0466064 (57 #5947) (reviewed by P. Damey; reprinted in Reviews in Number Theory, volume 5A, R24-229) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0351.12001 (Autorreferat) [scanned image].

[43] O. Taussky : Norms from quadratic gields and their relation to noncommuting 2 × 2 matrices III. A link between the 4-rank of the ideal class groups in ℚ(\sqrt{m}) and in ℚ(\sqrt{−m}), Math. Z. 154 (1977), Heft 2, 91–95; DOI: 10.1007/BF01241822 [h01845456vu48153.pdf (241.6 KB)];

                                                            PPN266833020_0154_log17.pdf (397,55 KB), GDZ; [PDF],

DigiZeitschriften; MR0506166 (58 #22014) (abstract) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0345.12007 (Autorreferat) [scanned image], Zbl 0345.12003 [scanned image].

[44] P. Morton : On Rédei’s theory of the Pell equation, J. Reine Angew. Math. 307/308 (1979), 373– 398; doi: 10.1515/crll.1979.307-308.373 [PDF (2633 KB)];

                                                            PPN243919689_0307_0308_log27.pdf (2,29 MB), GDZ;

[PDF], DigiZeitschriften; MR0534233 (81f:12005) (reviewed by W. C. Waterhouse; reprinted in Reviews in Number Theory, volume 5A, R14-316) [html, PDF, DVI, PostScript, Bib-TeX]; Zbl 0395.12018 (Autorreferat) [scanned image].

[45] J. C. Lagarias : On determining the 4-rank of the ideal class group of a quadratic field, J. Number Theory 12 (1980), no. 2, 191–196; doi:10.1016/0022-314X(80)90052-9 [PDF (305 K)]; MR0578811 (82c:12004) (reviewed by T. Nakahara; reprinted in Reviews in Number Theory, volume 5A, R14-322) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0438.12001 (reviewed by J. Stroeker) [scanned image].

[46] H. Kisilevsky : The Rédei-Reichardt theorem : Another proof, Ternary quadratic forms and norms (O. Taussky, ed.), Lecture Notes in Pure and Appl. Math., vol. 79, Dekker, New York and Basel, 1982, pp. 1–4; MR0437494 (55 #10422) (reviewed by Y. Furuta) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0512.12004 [scanned image].

[47] P. Morton : Density results for the 2-classgroups of imaginary quadratic fields, J. Reine Angew. Math. 332 (1982), 156–187; doi: 10.1515/crll.1982.332.156 [PDF (3502 KB)];

                                                            PPN243919689_0332_log10.pdf (2,87 MB), GDZ; [PDF], Di-

giZeitschriften; MR0656862 (83g:12004) (reviewed by E. Brown; reprinted in Reviews in Number Theory, volume 5A, R14-333) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0491.12004 (reviewed by K. S. Williams) [scanned image].

[48] P. Morton : Density results for the 2-classgroups of imaginary quadratic fields, Stud. Sci. Math. Hungar. 17 (1982), no. 1-4, 21–43; MR0761522 (85m:11067) (reviewed by C. J. Parry; reprinted in Reviews in Number Theory, volume 5B, R11-618) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0548.12001 (reviewed by K. S. Williams) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[49] F. Halter-Koch : Über den 4-Rank der Klassengruppe quadratischer Zahlkörper, J. Number Theory 19 (1984), no. 2, 219–227; doi:10.1016/0022-314X(84)90105-7 [PDF (417 K)]; MR0762771 (86a:11041) (reviewed by P. Damey; reprinted in Reviews in Number Theory, volume 5B, R29-651) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0543.12004 (reviewed by M. E. Keating) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[50] F. Gerth, III : The 4-class ranks of quadratic fields, Invent. Math. 77 (1984), fasc. 3, 489–515; DOI: 10.1007/BF01388835 [x66g82734j213u41.pdf (1.2 MB)];

                                                            PPN356556735_0077_log48.pdf (2,13 MB), GDZ; [PDF], Di-

giZeitschriften; MR0759260 (85j:11137) (reviewed by K. Komatsu; reprinted in Reviews in Number Theory, volume 5B, R11-615) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0533.12004 [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[51] F. Gerth, III and S. W. Graham : Application of a character sum estimate to a 2-class number density, J. Number Theory 19 (1984), no. 2, 239–247; doi:10.1016/0022-314X(84)90108-2 [PDF (344 K)]; MR0762774 (86b:11070) (reviewed by J. G. Hinz) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0551.12002 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[52] F. Gerth, III : The 4-class ranks of quadratic extensions of certain imaginary quadratic fields, Illinois J. Math. 33 (1989), Issue 1, 132–142; euclid.ijm/1255988809 [PDF (743 KB)]; MR0974015 (90b:11114) (reviewed by K. Girstmair; reprinted in Reviews in Number Theory, volume 5B, R29-757) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0663.12012 (reviewed by K. S. Williams) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[53] T. Uehara : On the 4-rank of the narrow ideal class group of a quadratic field, J. Number Theory 31 (1989), no. 2, 167–173; doi:10.1016/0022-314X(89)90017-6 [PDF (357 K)];MR0987569 (90e:11170) (reviewed by F. Diaz y Diaz; reprinted in Reviews in Number Theory, volume 5B, R29-766) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0681.12002 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[54] F. Gerth, III : The 4-class ranks of quadratic extensions of certain real quadratic fields, J. Number Theory 33 (1989), no. 1, 18–39; doi:10.1016/0022-314X(89)90057-7 [PDF (628 K)]; MR1014385 (90i:11125) (reviewed by A. Takaku; reprinted in Reviews in Number Theory, volume 5B, R29-773) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0694.12003 (reviewed by J. Hinz) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[55] P. Stevenhagen : Divisibility by 2-powers of certain quadratic class numbers, J. Number Theory 43 (1993), no. 1, 1– 19; doi:10.1006/jnth.1993.1001 [PDF (726 K)]; MR1200803 (94b:11113) (reviewed by K. S. Williams; reprinted in Reviews in Number Theory, volume 5B, R29-883) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0767.11054 (reviewed by R. Mollin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[56] P. Stevenhagen : The number of real quadratic number fields having units of negative norm, Experiment. Math. 2 (1993), No. 2, 121–136; euclid.em/1048516217 [PDF (297 KB), PostScript (NaN KB)]; stevenhagen.ps, ELibM, EMIS; MR1259426 (94k:11120) (reviewed by D. A. Buell; reprinted in Reviews in Number Theory, volume 5B, R11-765) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0792.11041 (reviewed by H. Cohn) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[57] 河野美文, 北村三郎, 中原徹 [Kohno, Y.; Kitamura, S.; Nakahara, T.] : グラフによる 2 次体の類群の 2 巾階数評価, 離散数理モデルにおける最適組合せ構造 [Optimal Combinatorial Structures on Discrete Mathematical Models] (京都大学数理解析研究所, 1992) (景山三平, 神保雅一編), 数理解析研究所講究録 [RIMS Kokyuroku] 820, 1993, 1–15; 0820-1.pdf; 0820-1.pdf (857.53 kB), URI: http://hdl.handle.net/2433/83181; MR1259324 (not reviewed).

[58] 河野美文, 中原徹 [Kohno, Y.; Nakahara, T.] : 2 次体の 2-類群構造の有向グラフ化について, 数理モデルの組合せ論的構造 [Combinatorial Structure in Mathematical Models] (京都大学数理解析研究所, 1993) (栗木進二編), 数理解析研究所講究録 [RIMS Kokyuroku] 853, 1993, 133–147; 0853-12.pdf; 0853-12.pdf (867.14 kB), URI: http://hdl.handle.net/2433/83735; MR1319200 (96b:11147) (reviewed by N. Kimura; reprinted in Reviews in Number Theory, volume 5B, R29-926) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0933.11501 (not reviewed).

[59] J. Hurrelbrink : Circulant graphs and 4-ranks of ideal class groups, Canad. J. Math. 46 (1994), no. 1, 169–183; DOI: 10.4153/CJM-1994-005-2 [cjm1994v46.0169-0183.pdf [1002 KB]]; MR1260342 (95h:11119) (reviewed by P. Stevenhagen; reprinted in Reviews in Number Theory, volume 5B, R29-910) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0792.05133 (summary) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[60] R. J. Kingan : Tournaments and ideal class groups, Canad. Math. Bull. 38 (1995), no. 3, 330–333; DOI: 10.4153/CMB-1995-048-1 [cmb1995v38.0330-0333.pdf [285 KB]]; MR1347306 (96f:05043) (reviewed by N. J. Pullman) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0877.11056 (reviewed by A. G. Earnest) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[61] A. Costa and F. Gerth, III : Densities for 4-ranks of totally complex quadratic extensions of real quadratic fields, J. Number Theory 54 (1995), no. 2, 274–286; doi:10.1006/jnth.1995.1118 [PDF (442 K)]; MR1354052 (96k:11124) (reviewed by P. Stevenhagen; reprinted in Reviews in Number Theory, volume 5B, R16-717) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0853.11089 (reviewed by C. Parry) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[62] W. Bosma and P. Stevenhagen : Density computations for real quadratic units, Math. Comp. 65 (1996), no. 215, 1327–1337; DOI: 10.1090/S0025-5718-96-00725-9 [PDF]; doi:10.2307/2153810 [PDF (275 KB)], JSTOR; MR1344607 (96j:11171) (reviewed by D. A. Buell; reprinted in Reviews in Number Theory, volume 5B, Y40-699) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0859.11064 (reviewed by H. Cohn) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[63] 末吉豊 [Sueyoshi, Y.] : 2 次体 Q(\sqrt{m}) と Q(\sqrt{−m}) の狭義 ideal 類群の 4-rank の比較, 代数的数論とフェルマーの問題 (京都, 1995) (小松啓一編), 数理解析研究所講究録 [RIMS Kokyuroku] 971, 1996, 134–

                                                            144; 0971-13.pdf; 0971-13.pdf (641.27 kB), URI:

http://hdl.handle.net/2433/60692; MR1481238 (98k:11160) (reviewed by H. Yokoi) [html, PDF, DVI, PostScript, Bib-TeX].

[64] F. Lemmermeyer : Rational quartic reciprocity II, Acta Arith. LXXX (1997), fasc. 3, 273–276; aa8036.pdf, icm; MR1451413 (98f:11004) (reviewed by L. Washington) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0901.11031 (reviewed by T. Metsänkylä) [html, PDF, XML, ASCII, DVI, PS, Bib-TeX].

[65] Y. Sueyoshi : On a comparison of the 4-rank of the narrow ideal class groups of Q(\sqrt{m}) and Q(\sqrt{−m}), Kyushu J. Math. [九州数学雑誌] 51 (1997), No. 2, 261–272; doi:10.2206/kyushujm.51.261, JOI:JST.JSTAGE/kyushujm/51.261 [51_261.pdf (742 K)]; 51_261.pdf (741 Kb), URL: http://hdl.handle.net/2324/ 11531; MR1470156 (98d:11130) (reviewed by F. Lemmermeyer) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0921.11055 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[66] Q. Yue [岳勤] : 2-Sylow subgroups of K_{2}O_{F} for real quadratic field and class group for imaginary quadratic field, Northeast. Math. J. [�北数学 (英文版)] 15 (1999), no. 3, 261–264; [CAJ, PDF], cnki; 中国数学文摘 m2000032943 [html]; MR1766667 (2001e:11113) (reviewed by J. Browkin) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1052.11509 (reviewed by O. Ninnemann) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[67] Q. Yue and K.-Q. Feng : The narrow 4-class rank of the tame kernel versus the 4-rank of the narrow ideal class groups of quadratic number fields, Acta Arith. XCVI (2000), fasc. 2, 155–165; doi:10.4064/aa96-2-4 [pdf_aa96-2-04.pdf (141.16 Kb)]; MR1814450 (2001m:11207) (reviewed by F. Lemmermeyer) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1052.11079 (reviewed by O. Ninnemann) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[68] Y. Sueyoshi : Relation between the narrow 4-class rank of quadratic number fields, Adv. Stud. Contemp. Math. (Pusan)

2 (2000), 47–58; MR1784626 (2001f:11187) (reviewed by F. Lemmermeyer) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0993.11060 (reviewed by E. Benjamin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[69] Y. Sueyoshi : On Rédei matrices with minimal rank, Far East J. Math. Sci. 3 (2001), no. 1, 121–128; MR1814957 (2001k:11225) (reviewed by F. Lemmermeyer) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0987.11068 (reviewed by T. Takeuchi) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[70] F. Gerth, III : Comparison of 4-class ranks of certain quadratic fields, Proc. Amer. Math. Soc. 129 (2001), no. 9, 2547–2552; DOI: 10.1090/S0002-9939-01-05922-6 [PDF]; doi:10.2307/2668777 [PDF (156 KB)], JSTOR; MR1838376 (2002c:11149) (reviewed by F. Lemmermeyer) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1012.11097 (summary) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[71] R. Osburn : A note on 4-rank densities, Canad. Math. Bull. 47 (2004), no. 3, 431–438; osburn8325.pdf [138 KB]; MR2072603 (2005f:11266) (reviewed by K. Hutchinson) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1127.11076 (reviewed by M. Kolster) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[72] M. Morishita [森下昌紀] : Ideal class groups and knot theory, 代数的数論とその周辺 [Algebraic number theory and related topics] (京都大学数理解析研究所, 2000) (栗原将人編), 数理解析研究所講究録 [RIMS Kokyuroku] 1376, 2004, 92–102; 1376-11.pdf; 1376-11.pdf (1.03 MB), URI: http://hdl.handle.net/2433/25621.

[73] T. Bülow Relative norms of units and 4-rank of class groups, Acta Arith. 119 (2005), fasc. 3, 243–264; doi:10.4064/aa119-3-2 [pdf_aa119-3-02.pdf (226.51 Kb)]; MR2167435 (2006f:11135) (reviewed by A. Mouhib) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1085.11052 (reviewed by C. Parry) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[74] É. Fouvry and J. Klüners : On the 4-class group of quadratic number fields, Invent. Math. 167 (2007), fasc. 3, 455–513; DOI: 10.1007/s00222-006-0021-2 [bu1k13m7753v6134.pdf (589.2 KB)]; 23819281.pdf (589 KB), Academic Search Premier, EBSCO; MR2276261 (2007k:11187) (reviewed by J. Browkin) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1126.11062 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[75] 末吉豊 : 2 次体の 4-rank の計算, 第 6 回北陸数論研究集会報告集 [Proceedings of the 6th Workshop on Number Theory] (金沢大学サテライトプラザ, 2007) (平林幹人, 野村明人, 山下浩, 木村巌, 菅野孝史編), 2008, pp. 1–21.

[76] 山村健 : 2 次 (拡大) 体の 4-rank に関する文献 (2008 年 4 月更新版), 第 6 回北陸数論研究集会報告集 [Proceedings of the 6th Workshop on Number Theory] (金沢大学サテライトプラザ, 2007) (平林幹人, 野村明人, 山下浩, 木村巌, 菅野孝史編), 2008, pp. 153–162.

[77] E. Fouvry and J. Klüners : On the Spiegelungssatz for the 4-rank, Algebra & Number Theory 4 (2010), no. 5, 493– 508; DOI: 10.2140/ant.2010.4.493 [ant-v4-n5-p01-s.pdf]; MR2679097 (2011i:11159) (reviewed by M. Horie) [html, PDF, DVI, PostScript, BibTeX]; Zbl pre05779833 (reviewed by B. Sodaïgui) [html, PDF, XML, ASCII, DVI, PS, Bib-TeX].

[78] É. Fouvry and J. Klüners : Weighted distribution of the 4-rank of class groups and applications, Int. Math. Res. Not. IMRN 2011, Issue 11, no. 16, 3618–3656; doi: 10.1093/imrn/rnq223 [3657.full.pdf]; MR2679097 (2011i:11159) (reviewed by M. Horie) [html, PDF, DVI, PostScript, BibTeX]; Zbl pre05779833 (reviewed by B. Sodaïgui) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[79] Q. Yue : Class groups under relative quadratic extensions, Acta Arith. 150 (2011), fasc. 4, 399–414; doi:10.4064/aa150-

                                                            4-5 [pdf_aa150-4-05.pdf (350.88 Kb)].

[80] L. Habsieger and E. Royer : Spiegelungssatz: a combinatorial proof for the 4-rank, Int. J. Number Theory 7 (2011), no. 8, 2157–2170; DOI: 10.1142/S1793042111005106 [PDF (222 KB)].

[81] F. Lemmermeyer : The 4-class group of real quadratic number fields, preprint, 1996.

Rédei の論文のハンガリー語版とドイツ語版との対応について:

[1] ⟷ [2], [5] (+[6]) ⟷ [8], [7] ⟷ [12],

[10] ⟷ [13], [11] ⟷ [14], [15] ⟷ [17],

[18] + [19] + [21] ⟷ [22].